Finding the Grade of a Mountain Trail. A straight trail with a uniform inclination leads from a hotel, elevation 5000 feet, to a lake in a valley, elevation 400 feet. The length of the trail is 4100 feet. What is the inclination (grade) of the trail?
Answer by stanbon(75887)

Finding the Grade of a Mountain Trail. A straight trail with a uniform inclination leads from a hotel, elevation 5000 feet, to a lake in a valley, elevation 400 feet. The length of the trail is 4100 feet. What is the inclination (grade) of the trail?
------------------------------------------------
grade = rise/run
---
grade = (5000-400)/4100
----
grade = 4600/4100
---
grade = 46/41 = 1.122
===========================
Cheers,
Stan H.